# Sim-to-Real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v2

@@ Introduction
Experimental setup. The training on the physical robot is conducted using a Husqvarna Research Platform (HRP) (MIT software licence) (Husqvarna 2017): a robotic lawnmower that we equip with an Nvidia Jetson AGX Orin development kit. The wheel dimensions in (1) are $r_{w}=12.25$ cm and $b=46.5$ cm. The training algorithms, based on the stable-baselines3 implementation (MIT license) (Raffin et al. 2021), are executed on the Jetson, with control signals sent to the HRP. The experimental environment is a $12\times 12$ meter indoor research arena. The agent receives its position from a high-precision Qualisys motion capture system (Qualisys 2023), which comprises 12 Oqus 700+ and 8 Arqus A12 cameras.
Simulation training details. We follow the training setup by Jonnarth, Zhao, and Felsberg (2024), and train the CPP policy in a simulated 2D environment. Specifically, we use the same environment parameters (also listed in the appendix), map geometries, and curriculum learning. We utilize soft actor-critic learning (Haarnoja et al. 2018b), and train for $2$-$8$M iterations with learning rate $2\cdot 10^{-5}$, batch size $256$, replay buffer size $5\cdot 10^{5}$, and discount factor $\gamma=0.99$. We use a $360^{\circ}$ simulated lidar field-of-view, $24$ lidar rays, $3.5$ m lidar range, $0.15$ m coverage radius, $0.26$ m/s maximum linear velocity, and $1$ rad/s maximum angular velocity. To improve the fidelity of the simulator we include a $0.5$ m/s2 maximum linear acceleration, $2$ rad/s2 maximum angular acceleration, $50$ ms action delay, and include the $10$ latest actions in the observation. The training time ranged from $25$ to $100$ hours on a T4 GPU and a 6226R CPU.
Real-world training details. For the fine-tuning, we lower the learning rate to $5\cdot 10^{-6}$ to avoid potential instability issues. In order not to give a disproportionately large weight to the first environment steps, we perform pure data collection during the first $5000$ steps, i.e. without any model updates. During real-world fine-tuning we use both fixed and randomized training maps that fit within the research arena, and set the goal coverage to $99\%$ for episode termination.
Evaluation. To evaluate the various RL policies, we measure the times $T_{90}$ and $T_{99}$ to reach $90\%$ and $99\%$ coverage, respectively. These metrics were chosen as the coverage time is usually important in CPP tasks, and common in the literature (Xu et al. 2022). To provide additional insights, we also measure path length and accumulated rotations, which are also of interest in some applications (Chen et al. 2019). We perform the evaluation on six maps not seen during training, see the appendix.
Extended analysis. We provide an extended analysis in the appendix, including videos, learned paths, collision statistics, performance on the individual evaluation maps, average speed and the learned entropy coefficient during fine-tuning, and an ablation study for the optimal training step size. The optimal training step size was $500$ ms, based on which, the state selection wait time in Table 1 was chosen. Finally, a step size of $500$ ms allows us to perform $4$ model updates for each environment step in the real setting.
Comparison of First-order CPP Policies
@@
We first investigate how well policies that assume a first-order Markov process transfer to our semi-virtual environment. In this experiment, the observation space does not include any previous actions, and the policy is trained for $8$M steps in a simulated environment without a limit on the linear and angular accelerations, and does not include any action delay. We evaluate the policy in simulation and in the real environment before and after fine-tuning. Since this policy does not depend on the previous action sequence, we can run it at any frequency. We evaluate it at the training frequency and as fast as possible, i.e. without waiting during state selection, which was measured to 15 ms per time step.
@@
By comparing coverage times between simulation and reality in Table 2, we see that they are in the same ball park, where the real results are slightly worse at the lower frequency. This is expected due to differences in kinematics and other sources of error. When increasing the inference frequency, we observe faster coverage times, even surpassing the simulation results. This is likely the case due to the fact that the agent can update its actions faster and quickly adapt to errors. Thus, it can navigate more efficiently and produce a smoother pattern.
@@
Table 2: First-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is 500 ms.
However, after fine-tuning, we find that the performance degrades. Our hypothesis is that as the observation includes no higher-order information, the optimal action cannot be deduced. For example, if the robot completely changes the direction of rotation, it takes some time before the new rotation takes effect. The observation may not change much, while the optimal action does. Thus, the performance degrades initially, which may take a long time to recover from. Running the inference at a high frequency partly circumvents this problem, explaining the high performance of the model without fine-tuning. While we did not evaluate the fine-tuned policy on a lower frequency, we expect it to perform worse as in the non fine-tuning case.
Figure 3: Average coverage times during fine-tuning on evaluation maps 1 and 2. Dashed line: No fine-tuning.
Figure 4: Average path length and accumulated full rotations during fine-tuning on evaluation maps 1 and 2.
Comparison of Higher-order CPP Policies
Next, we conduct a new experiment to evaluate how higher-order policies compare. We train a new model for $2$M steps in simulation, where we include the past 10 actions in the observation, and limit the maximum linear and angular accelerations based on realistic estimates. Furthermore, we include the measured real-world overhead as an action delay in simulation to better align the simulation with reality. Since the 10 actions correspond to specific points back in time, we keep the inference time step the same as during training.
The new model can learn higher-order Markovian dynamics, and in Table 3 we see that it transfers well when directly deployed from simulation, surpassing the first-order model for $99\%$ coverage under this lower inference frequency. Again, the real results are somewhat worse compared to the simulation results, as expected. In Figure 3, where we record the performance on a subset of the maps during fine-tuning, we observe that the performance degrades initially, but in contrast to the first-order model, the agent surpasses its original performance. $T_{99}$ degrades heavily in the early stage, while $T_{90}$ remains more or less constant. This suggests that the agent initially sacrifices long-term planning in favour of low-level control. After it has adapted to the low-level controls, then the long-term planning also improves. Comparing Tables 2 and 3, the higher-order policy surpasses the first-order policy in both $T_{90}$ and $T_{99}$ at $500$ ms after fine-tuning.
Furthermore, Figure 4 shows the path length and accumulated full rotations at $99\%$ coverage. The number of rotations were computed by summing the absolute differences in heading angle in every time step. Similar to the coverage time, both metrics degrade initially, and then improve over time, surpassing the initial performance. This shows that the agent finds a more efficient path after fine-tuning. It reduces the number of turns, which take time, while it overlaps its previous path less, thus reducing the total path length.
In total, the fine-tuning took roughly $8$ hours, which is only a small fraction of the weeks it would have taken to train from scratch. Note that, while training requires more computational power, such as a GPU, the computational requirements at inference time are much lower, and can even be done on a CPU (Jonnarth, Zhao, and Felsberg 2024).
@@
Table 3: Higher-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. $\Delta t$: Inference step size (ms). Note: Training step size is 500 ms.
Conclusions

### v3

@@ V-A Implementation Details
Real-world training details. In our real-world experiments, we evaluate our approach on the lawn mowing task. The fine-tuning on the physical robot is conducted using a Husqvarna Research Platform (HRP) (MIT software licence) [49]. This is a robotic lawnmower equipped with special firmware that allows it to be controlled through ROS. We equip the HRP with an Nvidia Jetson AGX Orin development kit. The wheel dimensions in (1) are $r_{w}=12.25$ cm and $b=46.5$ cm. The training algorithms, based on the Stable-Baselines3 implementation (MIT license) [45], are executed on the Jetson, with control signals sent to the HRP. The experimental environment is a $12\times 12$ meter indoor research arena. The agent receives its position from a high-precision Qualisys motion capture system [50], which comprises 12 Oqus 700+ and 8 Arqus A12 cameras. We use a $360^{\circ}$ simulated lidar field-of-view, $24$ lidar rays, $3.5$ m lidar range, $0.15$ m coverage radius, $0.26$ m/s maximum linear velocity, and $1$ rad/s maximum angular velocity. To improve the fidelity of the simulator during our sim-to-real experiments, we include a $0.5$ m/s2 maximum linear acceleration, $2$ rad/s2 maximum angular acceleration, $50$ ms action delay, and include the $10$ latest actions in the observation. For the real-world fine-tuning, we lower the learning rate to $5\cdot 10^{-6}$ to avoid potential instability issues. In order not to give a disproportionately large weight to the first environment steps, we perform pure data collection during the first $5000$ steps, i.e. without any model updates.
Environment details. Based on initial experiments, we find suitable reward parameters. The episodes are prematurely truncated if $\tau=1000$ consecutive steps have passed without the agent covering any new space. We set the maximum coverage reward $\lambda_{\mathrm{area}}=1$, the incremental TV reward scale $\lambda_{\mathrm{TV}}^{\mathrm{I}}=0.2$ for exploration and $\lambda_{\mathrm{TV}}^{\mathrm{I}}=1$ for lawn mowing, the collision reward $R_{\mathrm{coll}}=-10$, and the constant reward $R_{\mathrm{const}}=-0.1$. Note that $\lambda_{\mathrm{TV}}^{\mathrm{I}}$ is the only hyperparameter that differs between the two settings. The global TV reward scale was set to $\lambda_{\mathrm{TV}}^{\mathrm{G}}=0$ as it did not contribute to a performance gain in our ablations in Section V-B4. For the multi-scale maps, we use $m=4$ scales with $32\times 32$ pixel resolution, a scale factor of $s=4$, and $0.0375$ meters per pixel for the finest scale. Thus, the maps span a square with side length $76.8$ m. For more details regarding the choice of hyperparameters, see Appendix B. We apply curriculum learning, progressively increasing the environment complexity, and the goal coverage rate from $90\%$ to $99\%$ during training. For more details, see Appendix C. To increase the variation of encountered scenarios, we use both a fixed set of maps and procedurally generated maps, by randomizing grid-like floor plans. See Fig. 8 for examples, and Appendix D for more details. During real-world fine-tuning we use both fixed and randomized training maps that fit within the research arena, and set the goal coverage rate to $99\%$ for episode termination.
TABLE III: Time in seconds for reaching $90\%$ and $99\%$ coverage on Explore-Bench in simulation. Our method surpasses both frontier-based methods and a recent RL-based approach using active neural SLAM.
Map $\rightarrow$ Loop Corridor Corner Rooms Comb. 1 Comb. 2 Total
Method $\downarrow$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Distance frontier 124 145 162 169 210 426 159 210 169 175 230 537 1054 1662
RRT frontier 145 180 166 170 171 331 176 211 141 192 249 439 1048 1523
Potential field frontier 131 152 158 162 133 324 156 191 165 183 224 547 967 1559
Active Neural SLAM 190 214 160 266 324 381 270 315 249 297 588 755 1781 2228
Ours 89 101 98 128 80 291 82 99 87 93 120 349 556 1061
First- vs higher-order policies. In our experiments, we train two types of policies. For the first one, inertia and action delay are excluded from the simulation, i.e. the agent reaches the desired linear and angular velocities instantaneously, and the action takes effect immediately. Additionally, the observation space does not include any information about previous states. The action of this model is independent of the previous action sequence and only depends on the current state. Thus, we refer to it as a first-order policy. For the second type of policy, we consider inertia by limiting the maximum linear and angular accelerations, and also include an action delay. This captures higher-order Markovian dynamics, as the next state depends on a finite set of previous states and actions. Policies trained in this setting also observe the 10 previous actions, and we refer to them as higher-order policies.
Evaluation. To evaluate the various RL policies, we measure the times $T_{90}$ and $T_{99}$ to reach $90\%$ and $99\%$ coverage, respectively. We use these as the main metrics during ablations and comparisons with other methods, as the coverage time is usually important in CPP tasks, and common for benchmarking in the literature [6]. Some applications also value path length and accumulated rotations [25], so we also provide these. For further analysis, we also measure the collision frequency, the average speed, and the learned entropy in our fine-tuning experiments. The evaluation is performed on maps that are not seen during training. For details on training and evaluation maps used in simulation and in the real world, see Appendix F. Qualitative results in the form of learned paths can be found in Fig. 15, and videos can be found online. [^2] Videos can be found online at this link: https://drive.google.com/drive/folders/1J0vpOBuRhCHOxsnAE1Vd1cZ6-qHPlvUh?usp=sharing
V-B Simulation Experiments
Before going to a realistic setting, we perform comparisons and ablation studies in simulation in this section. Here, we train first-order policies unless stated otherwise.
TABLE IV: Time in minutes for reaching $90\%$ and $99\%$ coverage on the lawn mowing task in simulation.
Map $\rightarrow$ Map 1 Map 2 Map 3 Map 4 Map 5 Map 6 Total
Setting Method $\downarrow$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
Offline TSP 45 49 45 51 43 48 55 61 27 30 110 122 325 361
BSA 30 35 29 35 31 36 34 41 17 23 88 100 229 270
Online TSP 62 69 70 77 67 75 70 78 37 41 142 158 448 498
Ours 44 60 40 50 43 49 40 69 25 32 118 149 310 409
V-B1 Omnidirectional Exploration
In omnidirectional exploration, the agent observes its surroundings in all directions through a 360 field-of-view lidar sensor. For this setting, we evaluate our method on Explore-Bench [6], which is a recent benchmark that implements challenging environments, where four methods have been evaluated by the authors. These include three frontier-based methods, namely a distance-based frontier method [22], an RRT-based frontier method [23], and a potential field-based frontier method [24]. The fourth method is an RL-based approach, where Xu et al. [6] train an RL model to determine a global goal based on active neural SLAM [43]. The benchmark contains six environments; loop, corridor, corner, rooms, combination 1 (rooms with corridors), and combination 2 (complex rooms with tight spaces), which can be found in Appendix F. The results are presented in Table III. Our approach surpasses the performance of both the frontier-based methods and Active Neural SLAM. This shows that learning control signals end-to-end with RL is, in fact, a suitable approach to CPP. Furthermore, Fig. 1 (left) shows that the agent has learned an efficient exploration path in a complex and obstacle-cluttered environment.
V-B2 Non-Omnidirectional Exploration
We further evaluate our method on non-omnidirectional exploration, with a 180 lidar field-of-view. We reimplement and compare with the recent frontier-based RL method of Hu et al. [29], which was trained specifically for this setting. Their method uses RL to navigate to a chosen frontier node. The coverage over time is presented in Fig. 9. Our method outperforms the frontier-based RL approach, demonstrating that end-to-end learning of control signals is superior to a multi-stage approach for adapting to a specific sensor setup.
Figure 9: Simulation coverage over time in non-omnidirectional exploration.
V-B3 Lawn Mowing
For the lawn mowing task, we compare with the backtracking spiral algorithm (BSA) [21], which is a common benchmark for this CPP variation. Note however that BSA is an offline method and does not solve the mapping problem. As such, we do not expect our approach to outperform it in this comparison. Instead, we use it to see how close we are to a good solution. We further implement and compare with an offline and online version of a baseline that combines A* [51] with a traveling salesman problem (TSP) solver on nodes in a grid-like configuration. This has previously been proposed for the lawn mowing task in the offline setting for small environments [52]. To make it feasible for our larger environments, a heuristic was required for distant nodes, and periodic replanning was used for the online case, see Appendix G. Table IV shows $T_{90}$ and $T_{99}$ for six maps numbered 1-6, which can be found in Appendix F. Compared to offline BSA, our method takes $35\%$ and $51\%$ more time to reach $90\%$ and $99\%$ coverage respectively. This is an impressive result considering the challenge of simultaneously mapping the environment. Moreover, our approach outperforms the online version of the TSP-based solution, and even surpasses the offline version for $90\%$ coverage. The limiting factors for TSP are likely the grid discretization and suboptimal replanning, which lead to overlap, see Appendix G.
V-B4 Ablation Study
In Tables V, VI, and VII, as well as in Fig. 10, we explore the impact of different components of our approach via a series of ablations. Since some baselines in Table VI struggled to reach $90\%$ or $99\%$ coverage, we provide the coverage at fixed times, instead of $T_{90}$ and $T_{99}$.
TABLE V: Training step size comparison in ms. $T_{90}$ and $T_{99}$: Average time in minutes for reaching $90\%$ and $99\%$ coverage in simulation.
First-order Higher-order
$\Delta t$ $T_{90}$ $T_{99}$ $T_{90}$ $T_{99}$
150 7.1 12.7 7.6 13.7
250 6.4 9.3 6.4 10.1
500 6.0 9.1 5.8 8.5
1000 6.1 9.6 6.4 9.1
TABLE VI: Coverage ($\%$) at 1500 and 1000 seconds for mowing (Mow) and exploration (Exp) respectively, comparing agent architecture (NN), TV rewards, and frontier map observation ($M_{f}$).
Settings Coverage
$R_{\mathrm{TV}}^{I}$ $R_{\mathrm{TV}}^{G}$ $M_{f}$ NN Mow Exp
MLP 81.4 27.3
CNN 93.2 88.4
SGCNN 85.1 91.5
SGCNN 72.6 80.6
SGCNN 96.6 89.0
SGCNN 97.8 93.0
Optimal training step size. In Table V, we investigate the impact of the training step size in simulation, both for first-order and higher-order policies. These policies were evaluated on the real-world evaluation maps, see Appendix F. In both cases, $500$ ms seems to be the best choice, considering $90\%$ and $99\%$ coverage times. Thus, we choose $500$ ms for our experiments. If the step size is small, it may take longer to train in simulation since the computation time per step is the same. More steps are needed to complete the same number of episodes. In Table V, we used the same number of training steps, which meant that the number of episodes were lower for the smaller step sizes. Meanwhile, with a lower step size, future rewards are more strongly discounted, and thus, long-term planning might be limited. It is possible that this could be remedied by tuning hyperparameters, such as the discount factor, $\gamma$. We simply used the default value of $\gamma=0.99$ for SAC [15]. Finally, a lower step size requires a higher number of good consecutive actions to a achieve a good result, which might inhibit learning. On the other hand, if the step size is too large, the short-term control is limited, as each action is applied over a longer time period. This requires the agent to be more precise in its actions, which makes the problem harder. A step size of $500$ ms seems to strike a good balance. Based on this result, the state selection wait time in Table I was chosen. Finally, a step size of $500$ ms allows us to perform $4$ model updates for each environment step in the real setting.
TV rewards. In Table VI, we find that the incremental TV reward has a major impact on the lawn mowing problem, increasing the coverage from $85.1\%$ to $97.8\%$ (③ vs. ⑥). It also affects exploration, but not to the same extent. This is reasonable, as the total variation in the coverage map is lower in this case. When visualizing the learned paths, we found that by increasing $\lambda_{\mathrm{TV}}^{\mathrm{I}}$ beyond the optimal value with respect to coverage time, the agent learned more compact paths. For example, it would zig-zag rectangular regions parallel to the short edge instead of the long edge, and in general try to prioritize the nearest empty space. This behavior resulted in longer coverage times due to an increased number of turns. However, these paths could be considered more visually pleasing, and in turn be more appealing in some applications. Finally, as global TV has less temporal variation, and behaves more like a constant reward, it turns out not to be beneficial for CPP (⑤ vs. ⑥).
Agent architectures. The inductive prior of the CNN architecture enables a better understanding of the environment, and outperforms the MLP baseline (② vs. ①) in Table VI. Meanwhile, our proposed SGCNN architecture, which groups the different scales and convolves them separately, further improves the performance compared to a naive CNN (⑥ vs. ②), which treats all scales as the same spatial structure. We believe that the reason for the increased performance is that, with SGCNN, the agent is not faced with the additional challenge of discerning how the channels are spatially related. Thus, the representation learning becomes easier, and the agent can learn coverage paths more efficiently.
Multi-scale frontier maps. We further find in Table VI that our proposed frontier map representation is crucial for achieving a high performance with the multi-scale map approach (④ vs. ⑥). The frontier representation efficiently highlights the locations of non-visited regions, which we believe is especially useful for long-term planning. Without this input feature, too much information is lost in the coarser scales, thus hindering long-term planning.
Figure 10: Exploration coverage for different number of scales in simulation.
Number of scales. In Fig. 10, we compare the coverage for different number of scales on the exploration task. It shows that only using one or two scales is not sufficient, where at least three scales are required to represent a large enough environment to enable long-term planning. Meanwhile, increasing the number of scales even further does not hamper the performance, which shows that our approach can be applied to even larger environments than what have been considered in our experiments. The discrepancy was not as large for the lawn mowing problem, as it is more local in nature with a lower coverage radius.
Robustness to noise. As the real world is noisy, we evaluate how robust our method is to noise in Table VII. We apply Gaussian noise to the position, heading, and lidar measurements during both training and evaluation. Thus, both the maps and the sensor data perceived by the agent contain noise. We consider three levels of noise and present $T_{90}$ and $T_{99}$ for the lawn mowing task. The result shows that, even under high levels of noise, our method still functions well. Our approach surpasses the TSP solution under all three noise levels.
TABLE VII: Coverage time in minutes for different levels of noise on the lawn mowing task in simulation.
Noise standard deviation Time
Position Heading Lidar $T_{90}$ $T_{99}$
0.01 m 0.05 rad 0.05 m 310 409
0.02 m 0.1 rad 0.1 m 338 486
0.05 m 0.2 rad 0.2 m 317 434
V-B5 Collision Statistics
In most real-world applications, it is vital to minimize the collision frequency, as collisions can inflict harm on humans, animals, the environment, and the robot. To gain insights into the collision characteristics of our trained agents, we tracked the collisions during evaluation. For exploration, it varied between once every 100100-$1000$ seconds in simulation, which is roughly once every 5050-$500$ meters. For lawn mowing, it varied between once every 150150-$250$ seconds in simulation, which is roughly once every 3030-$50$ meters. These include all forms of collisions, even low-speed side collisions. The vast majority of the collisions were near-parallel, and we did not observe any head on collisions. The practical implications are very different between these cases. In the real setting, the collision frequency on the lawn mowing task on the six evaluation maps was measured to be once every $6.2$ minutes, or once every $65.4$ meters, on average. Again, the collisions were mainly from the side of the robot, and not head on.
V-C Real-world Experiments
In this section, we fine-tune our CPP agent in our semi-virtual environment, and evaluate how different policies generalize to the real world.
V-C1 Comparison of First-order CPP Policies
@@
We first investigate how well policies that assume a first-order Markov process transfer to our semi-virtual environment. In this experiment, the observation space does not include any previous actions, and the policy is trained for 88M steps in a simulated environment without a limit on the linear and angular accelerations, and does not include any action delay. We evaluate the policy in simulation and in the real environment before and after fine-tuning. Since this policy does not depend on the previous action sequence, we can run it at any frequency. We evaluate it at the training frequency and as fast as possible, i.e. without waiting during state selection, which was measured to 15 ms per time step.
@@
By comparing coverage times between simulation and reality in Table VIII, we see that the real results are slightly worse at the lower frequency. This is expected due to differences in kinematics and other sources of error. When increasing the inference frequency in the real world, we observe faster coverage times, even surpassing the simulation results. This is likely the case due to the fact that the agent can update its actions faster and quickly adapt to errors. Thus, it can navigate more efficiently and produce a smoother pattern. For the performance on the individual evaluation maps, see Appendix E.
However, after fine-tuning, we find that the performance degrades. Our hypothesis is that as the observation includes no higher-order information, the optimal action cannot be deduced. For example, if the robot completely changes the direction of rotation, it takes some time before the new rotation takes effect. The observation may not change much, while the optimal action does. Thus, the performance degrades, which may take a long time to recover from. This is problematic due to the lower training frequency. Running the inference at a high frequency partly circumvents this problem, explaining the high performance of the model without fine-tuning. While we did not evaluate the fine-tuned policy on a lower frequency, we expect it to perform worse as in the non fine-tuning case.
TABLE VIII: First-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. Inference step size is in ms. Note: Training step size is always 500 ms.
@@
TABLE IX: Higher-order policy. Average time in minutes for reaching $90\%$ and $99\%$ coverage on six evaluation maps. Env: Evaluation environment. Steps: Fine-tuning steps. Inference step size is in ms. Note: Training step size is always 500 ms.
@@
V-C2 Comparison of Higher-order CPP Policies
Next, we conduct a new experiment to evaluate how higher-order policies compare. We train a new model for 22M steps in simulation, where we include the past 10 actions in the observation, and limit the maximum linear and angular accelerations based on realistic estimates. Furthermore, we include the measured real-world overhead as an action delay in simulation to better align the simulation with reality. Since the 10 actions correspond to specific points back in time, we keep the inference time step the same as during training.
The new model can learn higher-order Markovian dynamics, and in Table IX we see that it transfers well when directly deployed from simulation, surpassing the first-order model for $99\%$ coverage under this lower inference frequency. Again, the real results are somewhat worse compared to the simulation results, as expected. Comparing Tables VIII and IX, the higher-order policy surpasses the first-order policy in both $T_{90}$ and $T_{99}$ at $500$ ms after fine-tuning. For the performance on the individual evaluation maps, see Appendix E.
Figure 11: Average coverage times during fine-tuning on two evaluation maps. Dashed line: No fine-tuning.
Figure 12: Average path length and accumulated full rotations during fine-tuning on two evaluation maps.
In Fig. 11, where we record the performance on a subset of the maps during fine-tuning, we observe that the performance degrades initially, but in contrast to the first-order model, the higher-order model surpasses its original performance. $T_{99}$ degrades heavily in the early stage, while $T_{90}$ remains more or less constant. This suggests that the agent initially sacrifices long-term planning in favour of low-level control. After it has adapted to the low-level controls, then the long-term planning also improves. Note that $T_{90}$ increases slightly after 40k iterations. It could be the case that the agent prioritizes the long-term goal over short-term performance towards the end of the fine-tuning, e.g. by choosing a less greedy path early on that reduces the final coverage time. Considering the general trend in the second half of Fig. 11, we expect the performance to continue to improve with further fine-tuning steps.
To further analyze the fine-tuned policy, we measure the path length, accumulated rotations, the average speed, and the learned entropy coefficient. Fig. 12 shows the path length and accumulated full rotations at $99\%$ coverage. The number of rotations were computed by summing the absolute differences in heading angle in every time step. Similar to the coverage time, both metrics degrade initially, and then improve over time, surpassing the initial performance. This shows that the agent finds a more efficient path after fine-tuning. It reduces the number of turns, which take time, while it overlaps its previous path less, thus reducing the total path length.
Figure 13: Average speed during fine-tuning on two evaluation maps. Dashed line: No fine-tuning.
Figure 14: The learned entropy coefficient in soft actor-critic [13] during fine-tuning.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e)
Figure 15: Learned paths for a fully trained agent, including the starting position (red triangle) and the end position (green square). Top row: Paths in simulation on the lawn mowing task (a-b) and on exploration (c-d). Bottom row: Paths in our semi-virtual environment on the lawn mowing task (a-d).
In Fig. 13, we analyse the average speed of the agent during fine-tuning. The initial drop in performance is also evident here. In the initial fine-tuning phase, the agent attains a lower average speed, as a result of e.g. more collisions and many small turns, while it adjusts to the new environment dynamics. Over time, it learns to control the robot in a more efficient manner, allowing it to keep a higher average speed with less turns, which results in lower coverage times.
Fig. 14 shows the learned entropy coefficient in the soft actor-critic framework [13]. It increases initially, suggesting that the agent prioritizes exploration over exploitation during the distribution shift to the real world. After the initial peak, the entropy coefficient decreases, putting more emphasis on exploitation. However, it does not reach the same level as in simulation, i.e. at 0 fine-tuning steps. This either suggests that the real environment contains more uncertainty, or that the learning process has not entirely converged yet. It is likely a combination of the two.
V-C3 Generalization
To evaluate the generalization capability of our approach, we perform a new exploration experiment, this time on a new robot platform. The new robot consists of an MiR100 platform integrated with a UR5e robotic arm (not used). The MiR platform has two driving wheels in the middle to control rotational and lateral movement, and four passive rotating wheels in each corner for stability. This allows for the linear and angular velocities to be controlled independently, and thus the predicted control signals can be passed directly to the robot platform, after proper renormalization. To further test the generalization capability, we change the sensor setup by severely limiting the visibility of the agent. We use the non-omnidirectional setup, and reduce the lidar field-of-view to $90^{\circ}$, i.e. $\pm 45^{\circ}$, and the range to $2$ meters. The inference is executed as fast as possible, i.e. with a step size of $15$ ms.
Figure 16: Coverage path on the MiR100 platform, including the start (red triangle) and end position (green square).
TABLE X: Coverage time in seconds in simulation and on the MiR100 platform on exploration with limited field-of-view ($90^{\circ}$) and range (22 meters). The mean and standard deviation is computed over five runs.
Environment Robot $T_{90}$ $T_{99}$
[…]
We perform the evaluation on a suitable map for this setting, which is shown in Fig. 16 together with a learned path. Table X compares $T_{90}$ and $T_{99}$ between simulation and the MiR platform. The real-world results are slightly worse, which can be expected when considering the discrepancy between the two settings. For example, the maximum linear and angular accelerations were lower on the MiR platform compared to the HRP, which was not accounted for in the simulation. However, the coverage results are within a reasonable margin, which demonstrates that our approach can indeed generalize to different sensor setups and robot platforms.
V-C4 Energy Efficiency
Due to the lightweight nature of our proposed SGCNN architecture, it is well suited for real-time applications. With only 800800k parameters, it is significantly smaller and faster than popular computer vision architectures, see Appendix A. Meanwhile, our full computational framework is energy efficient. When measuring the energy consumption on the battery powered Jetson module, it never exceeded 15 W during training, and drawing roughly 10 W during inference. Note that, while training requires more computational power, such as a GPU, the computational requirements at inference time are much lower, and can even be done on a CPU. In total, the fine-tuning took roughly $8$ hours, which is only a small fraction of the weeks it would have taken to train from scratch.
VI Conclusions
We present a method for online coverage path planning in unknown environments, based on a continuous end-to-end deep reinforcement learning approach. The agent implicitly solves three tasks simultaneously, namely mapping of the environment, planning a coverage path, and navigating the planned path while avoiding obstacles. For a scalable solution, we propose to use multi-scale maps, which is made viable through frontier maps that preserve the existence of non-covered space in coarse scales. We propose a novel reward term based on total variation, which significantly improves coverage time when applied locally, but not globally. The result is a task- and robot-agnostic CPP method that can be applied to different applications, which has been demonstrated on the exploration and lawn mowing tasks.